An airplane wing generates lift due to a differenc
Answer #1

Basic equation applicable = P + 0.5 \rhov^2

a) if both the wings are at same height, the velocity will same resulting in null pressure difference

b) Pbottom - Ptop = 0.5 \rho ( 245^2 - 222^2) = 0.5 ( 1.29 ) ( 60025 - 49284) =6927. 945 Pascals

c) Lift forc e= 6927. 945( 4) =27711.78 N

d) lift is directly proportional to dnsity, mars density is less than earth's so an airplane will experience less lift

e